# Explicitly Modeling Subcortical Vision with a Neuro-Inspired Front-End Improves CNN Robustness

### v1

@@ Abstract
Convolutional neural networks (CNNs) trained on object recognition achieve high task performance but continue to exhibit vulnerability under a range of visual perturbations and out-of-domain images, when compared with biological vision. Prior work has demonstrated that coupling a standard CNN with a front-end block (VOneBlock) that mimics the primate primary visual cortex (V1) can improve overall model robustness. Expanding on this, we introduce Early Vision Networks (EVNets), a new class of hybrid CNNs that combine the VOneBlock with a novel SubcorticalBlock, whose architecture draws from computational models in neuroscience and is parameterized to maximize alignment with subcortical responses reported across multiple experimental studies. Without being optimized to do so, the assembly of the SubcorticalBlock with the VOneBlock improved V1 alignment across most standard V1 benchmarks, and better modeled extra-classical receptive field phenomena. In addition, EVNets exhibit stronger emergent shape bias and overperform the base CNN architecture by 8.5% on an aggregate benchmark of robustness evaluations, including adversarial perturbations, common corruptions, and domain shifts. Finally, we show that EVNets can be further improved when paired with a state-of-the-art data augmentation technique, surpassing the performance of the isolated data augmentation approach by 7.3% on our robustness benchmark. This result reveals complementary benefits between changes in architecture to better mimic biology and training-based machine learning approaches.
@@ 1 Introduction
Convolutional neural networks (CNNs) have achieved remarkable performance across a range of object recognition benchmarks [NIPS2012_c399862d, simonyan2015deep, szegedy2014going, he2015deep, tan2020efficientnet], yet they remain vulnerable when faced with common corruptions [hendrycks2019benchmarking], domain shifts [geirhos2021partial, geirhos2022imagenettrained, geirhos2020accuracy], and adversarial perturbations [madry2019deep, goodfellow2015explaining]. These vulnerabilities not only limit deployment in real-world settings but also underscore fundamental disparities between computer vision models and primate vision [geirhos2022imagenettrained, geirhos2021partial]. In response, recent work has introduced biologically inspired models that integrate neuroscientific computations into CNN pipelines [Dapello2020.06.16.154542, cirincione2022implementing, pushpull2024, EVANS202296, 10.1007/978-3-031-44204-9_33]. A prominent example is the VOneNet family [Dapello2020.06.16.154542], which improves adversarial and corruption robustness by combining a biologically constrained front-end — the VOneBlock — to conventional CNN architectures. This block simulates processing in primate primary visual cortex (V1) via a fixed-weight, empirically constrained Gabor filter bank (GFB), nonlinearities reflecting responses of V1 simple and complex cells, and a neural noise generator. While VOneNets represent a key step towards neurally-aligned vision models, they abstract away the hierarchical processing in the early visual system, notably omitting subcortical circuits such as the retina and lateral geniculate nucleus (LGN). This raises the question of whether incorporating more complete models of V1, solely by refining upstream processing to include explicit subcortical processing, yields further gains in robustness and alignment with biological vision.
To address this question, we present the following key contributions:
@@
  • We introduce Early Vision Networks (EVNets), a new class of hybrid CNNs that combines two biologically-grounded modules — the VOneBlock and our new SubcorticalBlock — as a multi-stage front-end for a stardard CNN architecture.
  • We show that without any explicit optimization for V1 predictivity, EVNets improve neural and behavioral alignment with primate vision, outperforming both standard CNNs and VOneNets across multiple benchmarks. In particular, EVNets capture extra-classical RF phenomena more accurately, increase V1 predictivity and tuning property alignment, and better mimic human inductive biases through a stronger emergence of shape bias.
  • We demonstrate that EVNets deliver enhanced robustness across a diverse battery of evaluations, including adversarial attacks, common image corruptions, and domain shifts, while maintaining similar accuracy on clean images.
@@ 1.1 Related Work
The Difference-of-Gaussian (DoG) model emerged as the foundational linear framework for characterizing spatial summation over the receptive field (RF) of subcortical cells [kuffler1953, Rodieck1965-yv]. Subsequent extensions modeled extra-classical RF properties, including contrast gain control and surround suppression [Shapley1978-dx, Solomon2006-zt]. Building on this, divisive normalization [Carandini2011-xb, Bonin2005-cu] and cascading linear-nonlinear (LN) models [MANTE2008625] improved subcortical predictivity by incorporating the interaction between different visual processing stages. More recently, CNNs outperformed prior models in predicting subcortical responses to visual stimuli [McIntosh2016].
@@
CNN robustness improvements have largely been driven by data augmentation techniques [hendrycks2020augmix, hendrycks2021faces, 10.1007/978-3-031-19806-9_36]. To this end, standard benchmarks evaluate models under image corruptions [hendrycks2019benchmarking] and alternative renditions [hendrycks2021faces, salvador2022imagenetcartoon, wang2019learning, geirhos2021partial, geirhos2022imagenettrained]. Recent work highlights that composing augmentations [hendrycks2020augmix, hendrycks2021faces, 10.1007/978-3-030-58580-8_4], especially when integrated with architectural changes [pushpull2024], forwards the SOTA under this regime. Notably, PRIME augmentation [10.1007/978-3-031-19806-9_36] samples semantically-aligned transformations from maximum entropy distributions. In parallel, the vulnerability of CNNs to white-box adversarial perturbations has catalyzed extensive research [madry2019deep, obfuscated-gradients, carlini2019evaluating] with adversarial training emerging as the dominant paradigm for improving robustness [madry2019deep].
@@
A growing suite of metrics has emerged to quantify the alignment between models and the primate vision [Marques2021.03.01.433495, Freeman2013, SchrimpfKubilius2018BrainScore, Schrimpf2020integrative]. Metrics such as shape bias [geirhos2022imagenettrained] have been instrumental in measuring model-human behavioral alignment, while a parallel line of research emphasizes representational alignment through the comparison of model tuning properties to those observed in neural data [Marques2021.03.01.433495, Freeman2013, olaiya2022measuring]. Complementing these efforts, the BrainScore platform [SchrimpfKubilius2018BrainScore, Schrimpf2020integrative] provides a unified benchmark that integrates neural recordings and behavioral data across multiple visual cortical areas, including V1, V2, V4, and IT, alongside behavioral and task-driven metrics.
@@
Introducing biological computations into CNNs has consistently enhanced their robustness to input perturbations [EVANS202296, 10.1007/978-3-031-44204-9_33, pushpull2024, 10.1007/s00521-020-04751-8, Dapello2020.06.16.154542]. Gains have been observed with convolutional layers aligned to early visual RFs [EVANS202296, 10.1007/978-3-031-44204-9_33], push-pull inhibition motifs inspired by V1 [pushpull2024, 10.1007/s00521-020-04751-8], and by introducing the VOneBlock [Dapello2020.06.16.154542, cirincione2022implementing]. Besides improving adversarial and corruption robustness, VOneNets achieved improved accuracy in V1 predictivity, and, when combined with V1 divisive normalization, further sharpened alignment with V1 response properties and corruption robustness [cirincione2022implementing, olaiya2022measuring]. Finally, the systematic composition of hallmark V1 computations into CNNs recently achieved SOTA performance on explaining V1 predictivity and tuning properties [pogoncheff2023explaining].
@@ 2 Methods
Inspired by the hierarchical organization of early visual processing culminating in V1, we introduce EVNets, a new family of neuro-inspired CNNs that build upon the VOneNet framework. EVNets incorporate modular fixed-weight front-ends that reflects the functional stages of the early primate visual pathway (Fig. 1). This architecture comprises three key components: the SubcorticalBlock, modeling response characteristics of foveal neurons in the retina and the LGN; a variant of the VOneBlock, which models classical RF properties of V1 neurons; and a standard CNN back-end architecture. Together, the front-end blocks instantiate a composite cascading LN model of early cortical vision. EVNets adopt a spike-count formulation, abstracting over temporal dynamics, focusing solely on spatial encoding. The EVNet operates over a 7deg field-of-view (FoV), reduced from the 8deg used in the original VOneBlock [Dapello2020.06.16.154542]. We also extend the GFB to higher spatial frequencies (SFs), improving alignment with the SF tuning properties of primate V1 [DEVALOIS1982545] (cf. Supplementary Material A for an overview of VOneNets and our modifications). We trained three random seeds for both the VOneNet and the EVNet model families using a ResNet50 [he2015deep] back-end architecture (cf. Supplementary Material D.2 for a study on back-end generalization). All models were trained on the ImageNet-1k dataset [NIPS2012_c399862d], with clean accuracy evaluated on the standard validation split.
@@ 2.1 Architecture
where $r_{c}$ and $r_{s}$ are the center and surround radii, and $k_{s}/k_{c}$ is the peak contrast sensitivity ratio. We simulate biological color-opponent pathways characteristic of the different types of cells, with the P-cell stream incorporating red-green, green-red, and blue-yellow opponency, whereas M cells reflect achromaticity by incorporating a DoG filter with no color tuning [10.3389/fneur.2021.661938, Wiesel1966-zq].
@@
To mimic the subcortical mechanism of light adaptation [Carandini2011-xb, MANTE2008625, berardino2018eigendistortions], we introduce a biologically inspired module that performs global luminance normalization through a parameterized variant of the classical Naka-Rushton function [Naka1966-uo]. This transformation modulates the input $\mathbf{x}$ as
$\mathbf{x_{\text{LA}}}=\frac{\mathbf{x}^{\gamma}}{\mathbf{x}^{\gamma}+\bar{\mathbf{x}}^{\gamma}}-b_{\text{LA}}\text{,}$ (2)
where $x_{\text{LA}}$ is the light-adapted input, $\bar{\mathbf{x}}$ denotes the spatial average of the input, $\gamma$ controls the nonlinearity of the response, and $b_{\text{LA}}=1/2$ is a centering bias that ensures a null output for inputs matching the global mean.
@@
To model the adaptive gain control mechanisms characteristic of early visual processing [Carandini2011-xb, Bonin2005-cu, MANTE2008625, RaghavanENEURO.0515-22.2023, Mante2005-kc], we introduce a contrast normalization stage that normalizes activations by a local estimate of stimulus contrast. Formally, the normalized response is computed by
@@
The canonical push-pull pattern, emerging from the antagonistic interaction between ON- and OFF-center cells in early visual circuits, can be functionally approximated by subtracting rectified responses of opposite polarity pathways [doi:10.1126/science.8191289, Hirsch1998-lq]. Assuming ON and OFF cells of the same type share identical spatial profiles and gain, their RFs differ only in polarity. Under this assumption, since Equation 3 is antisymmetric with respect to $\mathbf{x}_{\text{DoG}}$, the contrast-normalized responses of the ON and OFF pathways, $\mathbf{x^{\pm}_{\text{CN}}}$, would satisfy $\mathbf{x^{+}_{\text{CN}}}=-\mathbf{x^{-}_{\text{CN}}}$. Thus, applying rectification to both signals and computing their differences gives
@@
Neuronal responses in the primate visual system exhibit trial-to-trial variability with distinct stochastic signatures across processing stages. In V1, spike count variability closely follows a Poisson distribution, where the variance is equal to the mean [Softky1993-te]. In this case, the Fano factor, defined as the ratio between the variance and the mean is equal to one. Conversely, subcortical neurons exhibit sub-Poisson variability, characterized by a spike count variance that grows slower than the mean [doi:10.1152/jn.00058.2022, doi:10.1152/jn.01171.2003]. To faithfully capture this hierarchical structure of neural noise, we implement a dual-source noise injection mechanism in which we add independent Gaussian noise to each units of both front-end blocks. This noise is calibrated to maintain an overall variability with a unit Fano factor, while trial-to-trial activations at the VOneBlock output exhibit heteroskedasticity consistent with V1 measurements [Softky1993-te]. Prior to noise injection, unit activations are linearly rescaled such that their mean response to a stimulus aligns with the empirically observed spike count of the corresponding primate neuronal population over a 50-ms integration window [doi:10.1152/jn.00058.2022, Dapello2020.06.16.154542].
@@ 2.2 Subcortical-Aligning Parameterization
Despite the wealth of empirical data on primate LGN and the existence of various fitted models [Carandini2011-xb, Bonin2005-cu, MANTE2008625, RaghavanENEURO.0515-22.2023, Mante2005-kc], the heterogeneity of modeling approaches across studies limits the direct reuse of parameters in the SubcorticalBlock, while complicating synthesis from the broader literature. To address this, we introduce a novel neurophysiologically-constrained hyperparameter tuning strategy designed to produce responses that best match the mean neuronal response properties of an LGN neuronal population taken from prior studies and modeling strategies. Specifically, we selected a total of $N=6$ different response property distributions measured at foveal LGN to ensure alignment in SF tuning, size tuning, and contrast sensitivity. The individual properties are: center, surround, excitation and inhibition radii [Solomon338, Rodieck1965-yv, Sceniak1999], suppression index [Solomon338] and saturation index [RaghavanENEURO.0515-22.2023].
We conducted a series of in silico experiments, presenting each cell with drifting gratings, quantifying each response property through the first harmonic (F1) of the cell’s response. We then performed hyperparameter search via Bayesian optimization [10.1007/3-540-07165-2_55], minimizing the loss
@@ 2.3 Model Evaluation
The emergence of shape bias was measured using a cue-conflict dataset [geirhos2022imagenettrained] that combines the shapes and textures of ImageNet samples using style transfer [7780634]. To quantify the correspondence between model internal representations and biological V1 responses, we assessed the activations from the first computational block of each model using two complementary BrainScore [SchrimpfKubilius2018BrainScore, Schrimpf2020integrative] benchmarks: V1 neural predictivity [Freeman2013] and V1 response property [Marques2021.03.01.433495]. V1 neural predictivity evaluates the degree to which model-derived features can account for the variance in primate V1 neuronal responses via partial least squares (PLS) regression. In contrast, V1 response property quantifies RF tuning similarity by comparing the distributions of 22 neural response properties — extracted from the same first-block activations — to empirical V1 distributions. These properties span 7 functional categories: orientation tuning, SF tuning, response selectivity, RF size, surround modulation, texture modulation, and response magnitude.
@@
We evaluated model corruption robustness by measuring top-1 accuracy on the ImageNet-C dataset [hendrycks2019benchmarking], which comprises 75 distinct corrupted variants of the ImageNet validation set. These corruptions are organized into 15 types applied at five severity levels, reflecting a specific real-world image degradation. The corruption types are further grouped into four broad categories: noise, blur, weather, and digital perturbations.
@@
To assess each model’s generalization under domain shifts, we averaged top-1 accuracies across five ImageNet-derived datasets that focus on renditions partially addressable by early visual processing mechanisms. Specifically, we used ImageNet-R [hendrycks2021faces], ImageNet-Cartoon [salvador2022imagenetcartoon], ImageNet-Drawing [salvador2022imagenetcartoon], ImageNet-Sketch [wang2019learning], and the 16-class Stylized-ImageNet [geirhos2021partial, geirhos2022imagenettrained] (Stylized16-ImageNet). Each dataset introduces representational changes such as abstraction, stylization, or domain-specific distortions while preserving the underlying semantic structure.
@@
Following the protocol in Dapello et al. [Dapello2020.06.16.154542], we evaluated robustness to white-box adversarial attacks by applying untargeted Projected Gradient Descent (PGD) [madry2019deep] on the same 5000 images of the ImageNet validation set. Attacks were carried out under $L_{\infty}$, $L_{2}$ and $L_{1}$ norm constraints and the perturbation budgets used were $\|\delta\|_{\infty}\in[1/1020,1/255,4/255,16/255]$, $\|\delta\|_{2}\in[0.15,0.6,1.2,2.4]$ and $\|\delta\|_{1}\in[40,160,640,2560]$. We used 16 PGD iterations with step size of $\|\delta\|_{p}/8$. Given the stochastic nature of our models, we employed the reparameterization trick [Kingma2014] during sampling to enable stable gradient-based learning. To account for the dual noise injection, we also performed Monte Carlo sampling to accurately estimate gradients under noise [obfuscated-gradients], averaging 32 samples per PGD iteration. Due to computational constraints, adversarial evaluations were conducted using two random seeds. Additional control experiments are provided in Supplementary Material C.1.
Ablation studies.
We performed a set of model ablations to assess the contribution of specific components to robustness. Details about these experiments are presented in Supplementary Material D.

### v2

@@ Abstract
Convolutional neural networks (CNNs) trained on object recognition achieve high task performance but continue to exhibit vulnerability under a range of visual perturbations and out-of-domain images, when compared with biological vision. Prior work has demonstrated that coupling a standard CNN with a front-end (VOneBlock) that mimics the primate primary visual cortex (V1) can improve overall model robustness. Expanding on this, we introduce Early Vision Networks (EVNets), a new class of hybrid CNNs that combine the VOneBlock with a novel SubcorticalBlock, whose architecture draws from computational models in neuroscience and is parameterized to maximize alignment with subcortical responses reported across multiple experimental studies. Without being optimized to do so, the assembly of the SubcorticalBlock with the VOneBlock improved V1 alignment across most standard V1 benchmarks, and better modeled extra-classical receptive field phenomena. In addition, EVNets exhibit stronger emergent shape bias and outperform the base CNN architecture by 9.3% on an aggregate benchmark of robustness evaluations, including adversarial perturbations, common corruptions, and domain shifts. Finally, we show that EVNets can be further improved when paired with a state-of-the-art data augmentation technique, surpassing the performance of the isolated data augmentation approach by 6.2% on our robustness benchmark. This result reveals complementary benefits between changes in architecture to better mimic biology and training-based machine learning approaches.  [^1] Code and model weights available at https://github.com/lucaspiper99/evnet/.
@@ 1 Introduction
Convolutional neural networks (CNNs) have achieved remarkable performance across a range of object recognition benchmarks [1, 2, 3, 4, 5], yet they remain vulnerable when faced with common corruptions [6], domain shifts [7, 8, 9], and adversarial perturbations [10, 11]. These vulnerabilities not only limit deployment in real-world settings but also underscore fundamental disparities between computer vision models and primate vision [8, 7]. In response, recent work has introduced biologically inspired models that integrate neuroscientific computations into CNN pipelines [12, 13, 14, 15, 16]. A prominent example is the VOneNet family [12], which improves adversarial and corruption robustness by combining a biologically constrained front-end — the VOneBlock — to conventional CNN architectures. This block simulates processing in primate primary visual cortex (V1) via a fixed-weight, empirically constrained Gabor filter bank (GFB), nonlinearities reflecting responses of V1 simple and complex cells, and a neural noise generator. While VOneNets represent a key step towards neurally-aligned vision models, they abstract away the hierarchical processing in the early visual system, notably omitting subcortical circuits such as the retina and lateral geniculate nucleus (LGN). This raises the question of whether refining upstream processing to V1, by explicitly modeling subcortical processing, yields further gains in robustness and alignment with biology. To address this question, we present the following key contributions:
@@
  • We introduce Early Vision Networks (EVNets), a new class of hybrid CNNs that combines two biologically-grounded modules, the VOneBlock with the new SubcorticalBlock, as a multi-stage front-end for a standard CNN architecture.
  • We show that without any explicit optimization for V1 predictivity, EVNets improve neural and behavioral alignment with primate vision, outperforming both standard CNNs and VOneNets across multiple benchmarks. In particular, EVNets capture extra-classical RF phenomena more accurately, increase V1 tuning property alignment, and better mimic human inductive biases through a stronger emergence of shape bias.
  • We demonstrate that EVNets deliver enhanced robustness across a diverse battery of evaluations, including adversarial attacks, common image corruptions, and domain shifts, and that these gains generalize to other architecture back-ends.
@@ 1.1 Related Work
The Difference-of-Gaussian (DoG) model emerged as the foundational linear framework for characterizing spatial summation over the receptive field (RF) of subcortical cells [17, 18]. Subsequent extensions modeled extra-classical RF properties, including contrast gain control and surround suppression [19, 20]. Building on this, divisive normalization [21, 22] and cascading linear-nonlinear (LN) models [23] improved subcortical predictivity by incorporating the interaction between different visual processing stages. More recently, CNNs outperformed prior models in predicting subcortical responses to visual stimuli [24].
Applications of subcortical vision.
Beyond modeling subcortical vision, DoG-based filtering emerged as a solution to edge detection [25], while Retinex theory [26] and lightness models [27] used spatial normalization to achieve color constancy. Subsequent frameworks such as scale-space representations [28] and multiscale Laplacian pyramids [29] generalized these computations into hierarchical contrast and boundary encoding. Recent work has reintroduced these ideas into deep architectures, by embedding explicit center–surround pathways for illumination-robust classification [30] and by unrolling Retinex-inspired optimization within CNNs for low-light image enhancement [31].
@@
CNN robustness improvements have largely been driven by data augmentation techniques [32, 33, 34]. To this end, standard benchmarks evaluate models under image corruptions [6] and alternative renditions [33, 35, 36, 7, 8]. Recent work highlights that composing augmentations [32, 33, 37], especially when integrated with architectural changes [14], forwards the SOTA under this regime. Notably, PRIME augmentation [34] samples semantically-aligned transformations from maximum entropy distributions. In parallel, the vulnerability of CNNs to white-box adversarial perturbations has catalyzed extensive research [10, 38, 39] with adversarial training emerging as the dominant paradigm for improving robustness [10].
@@
A growing suite of metrics has emerged to quantify the alignment between models and the primate vision [40, 41, 42, 43]. Metrics such as shape bias [8] have been instrumental in measuring model-human behavioral alignment, while a parallel line of research emphasizes representational alignment through the comparison of model tuning properties to those observed in neural data [40, 41, 44]. Complementing these efforts, the BrainScore platform [42, 43] provides a unified benchmark that integrates neural recordings and behavioral data across multiple visual cortical areas, including V1, V2, V4, and IT, alongside behavioral and task-driven metrics.
@@
Introducing biological computations into CNNs has consistently enhanced their robustness to input perturbations [15, 16, 14, 45, 12]. Gains have been observed with convolutional layers aligned to early visual RFs [15, 16], push-pull inhibition motifs inspired by V1 [14, 45], and by introducing the VOneBlock [12, 13]. Besides improving adversarial and corruption robustness, VOneNets achieved improved accuracy in V1 predictivity, and, when combined with V1 divisive normalization, further sharpened alignment with V1 response properties and corruption robustness [13, 44]. Finally, the systematic composition of hallmark V1 computations into CNNs recently achieved SOTA performance on explaining V1 predictivity and tuning properties [46].
@@ 2 Methods
Inspired by the hierarchical organization of early visual processing culminating in V1, we introduce EVNets, a new family of neuro-inspired CNNs that build upon the VOneNet framework. EVNets incorporate modular fixed-weight front-ends that reflects the functional stages of the early primate visual pathway (Fig. 1). This architecture comprises three key components: the SubcorticalBlock, modeling response characteristics of foveal neurons in the retina and the LGN; a variant of the VOneBlock, which models classical RF properties of V1 neurons; and a standard CNN back-end architecture. Together, the front-end blocks instantiate a composite cascading LN model of early cortical vision. EVNets adopt a spike-count formulation, abstracting over temporal dynamics, focusing solely on spatial encoding. The EVNet operates over a 7deg field-of-view (FoV), reduced from the 8deg used in the original VOneBlock [12]. We also extend the GFB to higher spatial frequencies (SFs), improving alignment with the SF tuning properties of primate V1 [47] (cf. Supplementary Material A for an overview of VOneNets and our modifications). We trained three random seeds for both the VOneNet and the EVNet model families using a ResNet50 [4] back-end architecture. All models were trained on the ImageNet-1k dataset [1], with clean accuracy evaluated on the standard validation split. Additional training details are provided in Supplementary Material E.4.
@@ 2.1 Architecture
where $r_{c}$ and $r_{s}$ are the center and surround radii, and $k_{s}/k_{c}$ is the peak contrast sensitivity ratio. We simulate biological color-opponent pathways characteristic of the different types of cells, with the P-cell stream incorporating red-green, green-red, and blue-yellow opponency, whereas M cells reflect achromaticity by incorporating a DoG filter with no color tuning [48, 49].
@@
To mimic the subcortical mechanism of light adaptation [21, 23, 50], we introduce a biologically inspired module that performs global luminance normalization. This transformation modulates the input $\mathbf{x}$ as
$\mathbf{x}_{\text{LA}}=\frac{\mathbf{x}-\bar{\mathbf{x}}}{\bar{\mathbf{x}}}\text{,}$ (2)
where $\mathbf{x}_{\text{LA}}$ is the light-adapted input and the $\bar{\mathbf{x}}$ denotes the average pixel intensity across channels and spatial dimensions of the input, ensuring a null output for pixel values matching the global mean.
@@
To model the adaptive gain control mechanisms characteristic of early visual processing [21, 22, 23, 51, 52], we introduce a contrast normalization stage that normalizes activations by a local estimate of stimulus contrast. The normalized response is computed by
@@
The canonical push-pull pattern, emerging from the antagonistic interaction between ON- and OFF-center cells in early visual circuits, can be functionally approximated by subtracting rectified responses of opposite polarity pathways [53, 54]. Assuming ON and OFF cells of the same type share identical spatial profiles and gain, their RFs differ only in polarity. Under this assumption, since Equation 3 is antisymmetric with respect to $\mathbf{x}_{\text{DoG}}$, the contrast-normalized responses of the ON and OFF pathways, $\mathbf{x^{\pm}_{\text{CN}}}$, would satisfy $\mathbf{x^{+}_{\text{CN}}}=-\mathbf{x^{-}_{\text{CN}}}$. Thus, applying rectification to both signals and computing their differences gives
@@
Neuronal responses in the primate visual system exhibit trial-to-trial variability with distinct stochastic signatures across processing stages. In V1, spike count variability closely follows a Poisson distribution, where the variance is equal to the mean, corresponding to a Fano Factor equal to one [55]. Conversely, subcortical neurons exhibit sub-Poisson variability, characterized by a spike count variance lower than the mean [56, 57]. To faithfully capture this hierarchical structure of neural noise, we implement a dual-source noise injection mechanism in which we add independent Gaussian noise to each unit of both front-end blocks scaled accordingly. This noise is calibrated at the unit level to maintain an overall variability with a unit Fano factor, while trial-to-trial activations at the VOneBlock output exhibit heteroskedasticity consistent with V1 measurements [55]. Prior to noise injection, unit activations are linearly rescaled such that their mean response to a stimulus aligns with the empirically observed spike count of the corresponding primate neuronal population over a 50-ms integration window [56, 12].
@@ 2.2 Subcortical-Aligning Parameterization
Despite the wealth of empirical data on primate LGN and the existence of various fitted models [21, 22, 23, 51, 52], the heterogeneity of modeling approaches across studies limits the direct reuse of parameters in the SubcorticalBlock, while complicating synthesis from the broader literature. To address this, we introduce a novel neurophysiologically-constrained hyperparameter tuning strategy designed to produce responses that best match the mean neuronal response properties of an LGN neuronal population taken from prior studies and modeling strategies. Specifically, we selected a total of $N=6$ different response property distributions measured at foveal LGN to ensure alignment in SF tuning, size tuning, and contrast sensitivity. The individual properties are: center, surround, excitation and inhibition radii [58, 18, 59], suppression index [58] and saturation index [51].
We conducted a series of in silico experiments, presenting each cell with drifting gratings, quantifying each response property through the first harmonic (F1) of the cell’s response. We then performed hyperparameter search via Bayesian optimization [60], minimizing the loss
@@ 2.3 Model Evaluation
Shape bias was measured using a cue-conflict dataset [8] that combines shapes and textures of ImageNet samples using style transfer [61]. To quantify the correspondence between model internal representations and V1 responses, we assessed the activations from the first block of each model using two complementary BrainScore [42, 43] benchmarks: V1 neural predictivity [41] and V1 response property [40]. V1 neural predictivity evaluates the degree to which model features can account for the variance in primate V1 responses via partial least squares (PLS) regression. In contrast, V1 response property quantifies RF tuning similarity by comparing the distributions of 22 response properties, extracted from the same first-block activations, to empirical V1 distributions. These properties span 7 functional categories: orientation and SF tuning, response selectivity, RF size, surround modulation, texture modulation, and response magnitude.
@@
We evaluated model corruption robustness by measuring top-1 accuracy on the ImageNet-C dataset [6], which comprises 75 distinct corrupted variants of the ImageNet validation set. These corruptions are organized into 15 types applied at five severity levels, reflecting a specific real-world image degradation. The corruption types are further grouped into four broad categories: noise, blur, weather, and digital perturbations.
@@
To assess each model’s generalization under domain shifts, we averaged top-1 accuracies across five ImageNet-derived datasets that focus on renditions partially addressable by early visual processing mechanisms. Specifically, we used ImageNet-R [33], ImageNet-Cartoon [35], ImageNet-Drawing [35], ImageNet-Sketch [36], and the 16-class Stylized-ImageNet [7, 8] (Stylized16-ImageNet). Each dataset introduces representational changes such as abstraction, stylization, or domain-specific distortions while preserving the underlying semantic structure.
@@
Following Dapello et al. [12], we evaluated robustness to white-box adversarial attacks by applying untargeted Projected Gradient Descent (PGD) [10] on 5000 images of the ImageNet validation set. Attacks were carried out under $L_{\infty}$, $L_{2}$ and $L_{1}$ norm constraints and the perturbation budgets used were $\|\delta\|_{\infty}\in[1/1020,1/255,4/255,16/255]$, $\|\delta\|_{2}\in[0.15,0.6,1.2,2.4]$ and $\|\delta\|_{1}\in[40,160,640,2560]$. We used 64 PGD iterations with step size of $\|\delta\|_{p}/32$. Additional implementation details are provided in Supplementary Material C.1.
EVNet variants.
We trained seven EVResNet50 variants derived from the full EVNet by performing six targeted ablations and one architectural addition (two seeds each). The ablations individually removed the P- and M-cell pathways, the contrast normalization and light adaptation layers, the VOneBlock, and the subcortical noise generator. When removing subcortical noise, cortical noise in the VOneBlock output was amplified to remain Poisson-distributed, consistent with the original VOneNet [12]. The final variant introduced an LGN–V2 skip connection [63] by concatenating the SubcorticalBlock output with the VOneNet bottleneck. While all EVResNet50 variants were tested for primate vision alignment and robustness, adversarial evaluations were restricted to a reduced attack set including only the two weakest perturbation strengths of each norm constraint.
Additional experiments.
To test whether improvements in adversarial robustness were not due to gradient masking [38], we performed a battery of controls according to the best practices [38, 64, 39] (cf. Supplementary Material C.1). We further evaluated the generalization of EVNet front-ends across back-end architectures by integrating them with EfficientNet-B0 [5] and CORnet-Z [65] (cf. Supplementary Material D.2). Finally, we assessed performance gains from multi-pass ensemble inference in Supplementary Material D.3.